OpenStudy (anonymous):

Use the given information to find the exact value of the expression. sin α =21/29 , α lies in quadrant II, and cos β =15/17 , β lies in quadrant I Find sin (α - β).

OpenStudy (anonymous):

Remember how to expand \(\sin(α - β)\)?

OpenStudy (anonymous):

@Calcmathlete , like this? sin(a-b) = s i n ( a - b )

OpenStudy (cwrw238):

lol

OpenStudy (anonymous):

lol. I thought you were going to hijack the question XD @sarah43 Remember the following? \[\sin(α - β) = \sinα\cosß - \cosα\sinß\]

OpenStudy (anonymous):

yes i did it but im not getting one of the answer choices

OpenStudy (anonymous):

What did you get and how did you get it?

OpenStudy (anonymous):

sin(21/29−15/17)=(sin21/29)(cos15/17)−(cos21/29)(sin15/17) @Calcmathlete

OpenStudy (campbell_st):

draw some triangles  find x and y using pythagoras... and then you can solve the problem y will be negative